# MARS: Meaning-Aware Response Scoring for Uncertainty Estimation in Generative LLMs

### v2

@@ Appendix B Conceptualizing the Response Semantics in Generative LLM Probabilities
Which planet is known as Red Planet?
It is Mars
@@
What is the capital city of Japan?
Tokyo is the capital city of Japan
@@
Which element has the chemical symbol "O"?
The chemical symbol "O" represents Oxygen
@@
Appendix C Training of BERT-like Model for Importance Function
As described in Section 3.3, we optimize the computational efficiency of MARS by training a single Bert-like model with 110M parameters to execute the importance function. This model is an adaptation of the pre-trained Bert-base-uncased [^5] https://huggingface.co/bert-base-uncased, modified by removing its last layer and incorporating two independent fully-connected (FC) layers. The first FC layer focuses on phrase detection with two output logits: “Begin Phrase” (BP) and “Inside Phrase” (IP), and classifies each token as BP if it marks the start of a phrase or as IP otherwise. This setup enables sentence segmentation into phrases. The second FC layer, tasked with assigning importance coefficients, produces a single output logit for each token’s importance coefficient.
For training data, we take a subset of 69192 question samples from the TriviaQA training set and questions of the whole training set of NaturalQA consisting of 87925. Then, we use these questions as input and feed them to all 7B-sized baseline models (Llama2-7b, Llama2-7b-chat, Mistral-7b, Falcon-7b) to yield the responses. This provides us with question-answer pairs. We use the Flair phrase chunking model to determine phrase labels in the answers, as described in Appendix C.1. For importance coefficient labels per token in the responses, we follow Algorithm  1.
@@ Appendix C Training of BERT-like Model for Importance Function
C.1 Dividing a Sentence to Phrases
To divide a sentence into phrases, we use the Flair phrase chunking model [^6] https://huggingface.co/flair/chunk-english Akbik et al. (2018), that uses 10 tags which are adjectival, adverbial, conjunction, interjection, list marker, noun phrase, prepositional, particle, subordinate clause and verb phrase. For example, the Flair model divides the sentence “The happy man has been eating at the dinner” as “The happy man”, “has been eating”, “at”, “the diner”.
C.2 Pseudocode of the Importance Function Algorithm
@@
TriviaQA
Which American-born Sinclair won the Nobel Prize for Literature in 1930?
Sinclair Lewis
Which musical featured the song Thank Heaven for Little Girls?
Gigi
What was the first movie western called?
Kit Carson
NaturalQA
When did the eagles win last super bowl?
2017
Who was the ruler of england in 1616?
James I
What is the hot coffee mod in san andreas?
a normally inaccessible mini-game
WebQA
what character did natalie portman play in star wars?
Padmé Amidala
what country is the grand bahama island in?
Bahamas
where did saki live?
United Kingdom
@@
Appendix D Experimental Details

### v3

@@ A.1 Discussion of the Differences with TokenSAR
Which planet is known as Red Planet? It is Mars
@@
What is the capital city of Japan? Tokyo is the capital city of Japan
@@
Which element has the chemical symbol "O"? The chemical symbol "O" represents Oxygen
@@
Appendix B Training of BERT-like Model for Importance Function
As described in Section 3.3, we optimize the computational efficiency of MARS by training a single Bert-like model with 110M parameters to execute the importance function. This model is an adaptation of the pre-trained Bert-base-uncasedhttps://huggingface.co/bert-base-uncased, modified by removing its last layer and incorporating two independent fully-connected (FC) layers. The first FC layer focuses on phrase detection with two output logits: “Begin Phrase” (BP) and “Inside Phrase” (IP), and classifies each token as BP if it marks the start of a phrase or as IP otherwise. This setup enables sentence segmentation into phrases. The second FC layer, tasked with assigning importance coefficients, produces a single output logit for each token’s importance coefficient.
For training data, we take a subset of 69192 question samples from the TriviaQA training set and questions of the whole training set of NaturalQA consisting of 87925. Then, we use these questions as input and feed them to all 7B-sized baseline models (Llama2-7b, Llama2-7b-chat, Mistral-7b, Falcon-7b) to yield the responses. This provides us with question-answer pairs. We use the Flair phrase chunking model to determine phrase labels in the answers, as described in Appendix B.1. For importance coefficient labels per token in the responses, we follow Algorithm  1.
@@ Appendix B Training of BERT-like Model for Importance Function
B.1 Dividing a Sentence to Phrases
To divide a sentence into phrases, we use the Flair phrase chunking modelhttps://huggingface.co/flair/chunk-english Akbik et al. (2018), that uses 10 tags which are adjectival, adverbial, conjunction, interjection, list marker, noun phrase, prepositional, particle, subordinate clause and verb phrase. For example, the Flair model divides the sentence “The happy man has been eating at the dinner” as “The happy man”, “has been eating”, “at”, “the diner”.
B.2 Pseudocode of the Importance Function Algorithm
@@
TriviaQA Which American-born Sinclair won the Nobel Prize for Literature in 1930? Sinclair Lewis
Which musical featured the song Thank Heaven for Little Girls? Gigi
What was the first movie western called? Kit Carson
NaturalQA When did the eagles win last super bowl? 2017
Who was the ruler of england in 1616? James I
What is the hot coffee mod in san andreas? a normally inaccessible mini-game
WebQA what character did natalie portman play in star wars? Padmé Amidala
what country is the grand bahama island in? Bahamas
where did saki live? United Kingdom
@@
Appendix C Experimental Details